OpenStudy (anonymous):

A batter hits a baseball upward with an initial speed of 96 feet per second. After how many seconds does the ball hit the ground? Use the formula h = rt - 16t2 where h represents height in feet and r represents the initial speed.

OpenStudy (bahrom7893):

after how many seconds does the ball hit the ground? that means h = 0 rt - 16t^2 = 0 r - initial speed

OpenStudy (bahrom7893):

So: 96t - 16t^2 = 0 16(6t - t^2) = 0

OpenStudy (bahrom7893):

6t-t^2 = 0 t(6-t)=0 t = 0 or t = 6

OpenStudy (anonymous):

Is he in the major leagues or minors.

OpenStudy (bahrom7893):

t = 6

OpenStudy (bahrom7893):

is the answer.. since t = 0 is initial value

OpenStudy (anonymous):

Well after at least 0 seconds, because it is impossible to have a negative time.. unless you go with time travel o.O

OpenStudy (bahrom7893):

t = 6 is the answer i meant

OpenStudy (bahrom7893):

guys random fun facts can go in chat.

OpenStudy (anonymous):

not funny^

OpenStudy (anonymous):

sorry bahrom, I just had to do it

OpenStudy (anonymous):

thanks bahrom7893 :) and very funny you guys

OpenStudy (anonymous):

interesting factorial mancake

OpenStudy (anonymous):

she thought it was funny

OpenStudy (anonymous):

I like dogs with mustard..